The demand for product X depends on the price of product X as well as the average household income (Y) according to the following relationship

Qdx = 550 - 10 P + 0.001Y

The supply of product X is positively related to own price of product X and negatively dependent upon W, the price of some input. This relationship is expressed as:

Qsx = 90 + 50 P - 6 W
Given that Y = 40,000 and W = 2, what is the:

1. Equilibrium price?
2. Equilibrium quantity?
1
Expert's answer
2020-12-11T11:11:26-0500

1) at equilibrium price

"Q_{dx}=Q_{sx}"

"550-10P\n+0.001Y=90+50P-6W"

"550-10P+0.001\times40000=90+50P-6\times2"

"550-10P+40=90+50P-12"

"550+40-90+12=50P+10P"

"512=60P"

"P=8.5333"

2) equilibrium quantity = "550-10\times 8.5333+0.001\times 40000=504.667"